OpenStudy (anonymous):

Estimate the circumference. use 3.14 or 22/7 to approximate "tt" ill give answers after yall figure it out :) medal for reward!!

OpenStudy (anonymous):

plz help....@TheЅmartOne

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

don't give direct answers..plz explain

OpenStudy (anonymous):

that is not a answer... A.220 ft B.440 ft C.879 ft D.1758 -@TheSmartOne

Elsa213 (elsa213):

@TheSmartOne *

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @Elsa213 c=2 pie r C= 2*22/7*7 c= 44 \(\color{blue}{\text{End of Quote}}\) It should be C = 2 * (22/7) * 140

sammixboo (sammixboo):

Circumference = 2 x pi x radius We know what our radius is, 140. We should use 3.14 for pi, so what is 2 x 3.14 x 140

Elsa213 (elsa213):

Thank you for making me look stupid, Smarty. >.<

sammixboo (sammixboo):

(make sure to round to get your answer)

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

anyone^ is it c?

Elsa213 (elsa213):

@sammixboo

TheSmartOne (thesmartone):

correct

OpenStudy (anonymous):

ok thx